OpenStudy (sparklestaraa):

A student needs to make a circular cardboard piece with an area between 154 square inches and 616 square inches. The function f(r) relates the area of the cardboard piece, in square inches, to the radius r in inches. Which of the following shows a reasonable domain for f(r)?

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (sparklestaraa):

B ?

OpenStudy (sparklestaraa):

@pooja195

OpenStudy (danjs):

\[154 \le \pi *r^2 \le 616\] the domain of the function f(r), is all the values for r

OpenStudy (danjs):

you need to solve that inequality for r

OpenStudy (danjs):

sorry, it says between, it is just 154 < pi*r^2 < 616 not equal to also